# Easy conditioning far beyond Gaussian

### v3

@@ Abstract
Estimating and sampling from conditional densities plays a critical role in statistics and data science, with a plethora of applications. Numerous methods are available ranging from simple fitting approaches to sophisticated machine learning algorithms. However, selecting from among these often involves a trade-off between conflicting objectives of efficiency, flexibility and interpretability. Starting from well known easy conditioning results in the Gaussian case, we show, thanks to results pertaining to stability by mixing and marginal transformations, that the latter carry over far beyond the Gaussian case. This enables us to flexibly model multivariate data by accommodating broad classes of multi-modal dependence structures and marginal distributions, while enjoying fast conditioning of fitted joint distributions. In applications, we primarily focus on conditioning via Gaussian versus Gaussian mixture copula models, comparing different fitting implementations for the latter. Numerical experiments with simulated and real data demonstrate the relevance of the approach for conditional sampling, evaluated using multivariate scoring rules.
Keywords: Stability by conditioning, Mixtures, Copula, Conditional sampling
@@ 1 Introduction
Modeling complex dependencies between variables is a fundamental task in statistics. Traditional methods often rely on Gaussian assumptions due to their mathematical tractability and ease of implementation. However, real-world data frequently exhibit features such as multi-modality, heavy tails, and skewness, which are not well-captured by Gaussian distributions. As a consequence, there is growing interest in more flexible approaches that can accommodate these characteristics while still permitting efficient inference and modeling.
In many applications, a comprehensive description of the stochastic behavior of some (multivariate) response variable given specific covariate values is required. Assuming a continuous framework, estimating a full conditional density allows computing a wide range of quantities of interest, such as confidence intervals or conditional quantiles. However, Conditional density estimation (CDE) is a challenging task since conditional distributions can be arbitrarily complex and subject to intricate changes throughout the input space.
Many methods have been proposed in the literature to perform the task of estimating conditional densities. Non-parametric methods for CDE such as conditional kernel density estimation [5] have been widely used but are susceptible to high-dimensional problems [14] and require careful bandwidth selection. Semi-parametric methods such as Least Square Density Ratio Estimation [25] have shown promising results in moderate dimensional problems. More flexible techniques, such as Distributional Random Forests [9], have demonstrated efficiency but they lack of explainability.
Parametric methods based on deep neural networks have also been used extensively for CDE. Mixture density networks [7] and kernel density networks [1], assume a mixture form for the conditional distribution. Normalising Flows [28] or Neural auto-regressive estimators [20] can also be used for this task. Recently, some modified versions of deep generative models allowing for conditioning such as conditional Variational Auto-Encoders (CVAE) [24], or conditional Generative Adversarial Networks (CGAN) [18] have shown efficiency for various tasks including image generation. These methods can learn complex distributions in high dimensions but require large amounts of training data and may be considered to lack interpretability.
@@
In this work, we are investigating and presenting results on a class of generative models rooted in multivariate statistics and enjoying an elegant theoretical construction. The goal is to estimate the conditional distribution of a random vector $\mathbf{X_{1}}\in\mathbb{R}^{q}$ given an observation $\mathbf{x_{2}}$ of $\mathbf{X_{2}}\in\mathbb{R}^{p}$, based on a sample of moderate size and taking values in a continuous range, $(\mathbf{X_{1}},\mathbf{X_{2}})$ previously acquired. This setting is common in statistics, such as in probabilistic forecasting or imputation of systematically missing values, for example in medical or epidemiological applications. We aim to define an approach for conditional density estimation that is efficient, flexible, interpretable, and privacy-preserving, while also not requiring extensive training data.
Assuming conditions of existence to be fulfilled, a natural approach to estimate the conditional density $f(\mathbf{x_{1}}|\mathbf{x_{2}})$ is to first estimate the joint density by $\hat{f}(\mathbf{x_{1}},\mathbf{x_{2}})$ and then perform conditioning. Most methods have traditionally focused on the Gaussian case, leveraging the availability of analytical formulas for conditional distributions. Gaussian copula regression [21], while allowing for arbitrary continuous marginal distributions, still assumes a Gaussian dependence structure, which provides the advantage of closed-form expressions for the conditional distributions and enables rapid conditional sampling [19]. However, this reliance on a Gaussian copula imposes restrictive assumptions, particularly in its inability to capture complex (tail) dependencies and asymmetries commonly found in observational data.
By introducing the concept of trans-dimensional families of probability distributions stable by conditioning and demonstrating that this concept extends to mixtures and transformations, we show that a much broader class of probability distributions benefits from analytical conditioning. For applications, we will mainly focus on the Gaussian mixture copula model (GMCM) [27] due to its ease of implementation and its ability to capture multi-modal dependencies. Additionally, we demonstrate that incorporating automatic differentiation enhances the parameter fitting process of GMCM compared to conventional approaches for computing derivatives [26]. It is worth noting that this model can also be employed in a purely generative framework, such as for the privacy-preserving generation of synthetic data.
@@
As a motivating example, we will consider the Wine data-set, which represents chemical analysis of wines grown in the same region in Italy. The analysis determined the quantities of 13 constituents found in each of the three types of wines. For illustrative purposes, we will focus on $2$ constituents, the Alcohol level and Malic acids concentration. Figure 1 illustrates the approach proposed using these variables. The joint density appears here to be rather well modeled by a GMCM, which conditional distribution can be calculated and simulated from efficiently as detailed throughout the paper.
@@
Figure 1: Different steps for estimating a conditional density: from a sample of the joint distribution (left), we estimate the joint density by a GMCM (middle) and for a new value of Malics acids concentration, we estimate (or sample from) the conditional distribution of Alcohol level by using analytical formulas (right).
@@
The remainder of the paper is organized as follows. In Section 2, we introduce the property of stability by marginalization and conditioning for families of multivariate distributions, provide some examples of families that satisfy these properties and show that it can be extended to mixtures and transformations. In Section 3, we present our method for estimating efficiently a conditional distribution by assuming a parametric form for the joint distribution and leveraging analytical formulas for conditioning. Section 4 focuses on the implementation of our algorithm for a particular case which is the GMCM. For the demonstration of the usefulness of our method, we perform numerical simulations and include real data examples in Section 5. Finally, in Section 6, we draw conclusions and outline areas for future research.
@@ 2.1.1 Definition
Let us first consider a fixed dimension $d\geq 1$ and a product of spaces $\prod_{k=1}^{d}\mathcal{X}_{k}$. The $\mathcal{X}_{k}$ are assumed to be measure spaces, typically $\mathbb{R}$ or its subspaces, each endowed with the Lebesgue measure or its restrictions. We consider absolutely continuous probability distributions on $\prod_{k=1}^{d}\mathcal{X}_{k}$ with positive probability density functions $f^{(d)}(\cdot;\theta)$, where $\theta$ denotes a parameter. In this framework, we work indifferently with probability density functions (pdfs) or cumulative distribution functions (cdfs).
@@
Let us now introduce sets of indices that will be used for marginal distributions. For a given dimension $d\geq 1$, the sets of $\ell$ strictly ordered indices are:
@@
  • $\mathbf{i}=(i_{1},\ldots,i_{\ell})$ set of $\ell$ indices with $1\leq i_{1}<\ldots<i_{\ell}\leq d$, $11\leq\ell\leq d-$ 1,
@@
For $f$ a positive pdf on $\prod_{k=1}^{d}\mathcal{X}_{k}$ and a set of $\ell$ strictly ordered indices $\mathbf{i}\in I_{\ell}^{(d)}$, we define the function:
[…]
which applies the marginalization of a $d$-dimensional density over $\ell$ variables.
To extend this operation to a trans-dimensional setting, we extend the marginalization operator as follows. For $d\geq 1$ and $f^{(d)}(\cdot;\theta)\in\mathcal{F}^{(d)}$, we define :
@@
A trans-dimensional family of probability distributions $\mathcal{F}$ is said to be stable by marginalization if $\text{Marg}(\mathcal{F})\subset\mathcal{F}$.
Let us know consider conditioning. Since the pdfs considered here are positive, all conditional distributions are defined. For $f$ a positive pdf on $\prod_{k=1}^{d}\mathcal{X}_{k}$, a pair of sets of strictly ordered indices $(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d}$ and observed components $(x_{j_{1}},\ldots,x_{j_{m}})$, we define:
@@
In the same manner as above for marginalization, we define a conditioning operator taking a pdf in a parametric family as input and returning every possible conditional densities i.e for all pairs of sets of strictly ordered indices and every possible set of conditioning variables. For a function $f^{(d)}(\cdot;\theta)\in\mathcal{F}^{(d)}$, we have :
@@
A trans-dimensional family of probability distributions $\mathcal{F}$ is said to be stable by conditioning if $\text{Cond}(\mathcal{F})\subset\mathcal{F}$.
@@ 2.1.2 Examples
Stability by marginalization and conditioning can be demonstrated by showing that the marginal or conditional distributions remain within the same family of distributions, albeit in a lower dimension. Here, we will present several examples of distribution families that exhibit this property.
@@
    where $\Theta^{(d)}=\mathbb{R}^{d}\times\text{PD}_{d}$ where $\text{PD}_{d}$ is the set of all symmetric positive definite matrix on $\mathbb{R}^{d\times d}$. The trans-dimensional family of Gaussian distributions is defined by: $\mathcal{F}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}$ with $D=+\infty$. By properties of the multivariate Gaussian distribution, we know that the marginal and conditional distributions remains Gaussian with parameters that can be computed analytically via well-known formulas. We have that $\text{Marg}(\mathcal{F})\subset\mathcal{F}$ and $\text{Cond}(\mathcal{F})\subset\mathcal{F}$, the trans-dimensional family of Gaussian distributions is stable by marginalization and by conditioning.
  • Student $t$-distribution: Another example is the family of multivariate Student t-distributions, which is a generalization for multivariate normal distributions allowing for heavier tails. Unlike the multivariate Gaussian distribution, it can model tail dependencies and it includes an additional “degrees of freedom” parameter, which controls the shape of the distribution, with higher values leading to distributions that resemble the multivariate Gaussian, while lower values result in heavier tails.
@@
    The multivariate Student $t$-distribution is stable by marginalization and conditioning [11]; more details are included in the appendix.
  • Other examples: More generally the family of elliptically contoured distributions also enjoys these properties [8]. Other examples of families of multivariate distributions stable by marginalization and conditioning are the multivariate unified skew normal and skew Student $t$-distributions [3]. Unlike elliptical distributions, they allow for skewness.
  • Counter example: Let us now consider a simple counter-example of a family of multivariate distributions which is stable by marginalization but not stable by conditioning. For a given $d\geq 1$, consider the family of multivariate Student $t$-distributions with a fixed degrees of freedom $\nu\in\left[0,+\infty\right[$ i.e
@@
    where $f^{(d)}(\mathbf{x};\mu,\Sigma)=\frac{\Gamma(\frac{\nu+d}{2})}{\Gamma(\frac{\nu}{2})(\nu\pi)^{\frac{d}{2}}|\Sigma|^{\frac{1}{2}}}(1+\nu^{-1}(\mathbf{x}-\mu)^{T}\Sigma^{-1}(\mathbf{x}-\mu))^{-\frac{(\nu+d)}{2}}$ and the trans-dimensional family is defined by $\mathcal{F}_{\nu}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}_{\nu}$ with $D=+\infty$. Then by properties of the multivariate Student $t$-distribution [11], this family is stable by marginalization but not by conditioning, because when conditioning, we change the degrees of freedom of the distribution.

### v4

@@ Abstract
We address the challenge of conditioning multivariate densities, extending analytical conditioning results far beyond the Gaussian case. We review and discuss families of multivariate distributions that do enjoy analytical conditioning, also providing a few counter-examples. Proving that trans-dimensional stability under conditioning extends to mixtures and transformations, we demonstrate that a broader class of multivariate distributions inherit easy conditioning properties. Building on this insight, we developed a generative method to estimate conditional distributions from data by first fitting a flexible joint distribution using copulas and then performing analytical conditioning in a latent space. We specifically apply this methodology to Gaussian Mixture Copula Models (GMCM) and examine various fitting strategies. Through simulations and real-world data experiments, we showcase the efficacy of our method in tasks involving conditional density estimation and data imputation.
@@ 1 Introduction
Conditioning a multivariate density is a crucial yet challenging task in statistics. While conditional densities are easily obtained in the Gaussian case, more flexible multivariate distributions often require computationally intensive methods to obtain approximate conditional samples. In contrast, fast (approximate) evaluations of conditional probability density functions and cumulative distributions functions may be advantageous for computing quantities of interest, such as conditional quantiles and dependence measures.
In this paper, we explore non-Gaussian families of multivariate distributions that allow analytical computation of conditional distributions such as multivariate Student $t$-distribution [ding2016conditional] or multivariate skew normal distributions [azzalini1996multivariate]. Furthermore, by introducing the concept of trans-dimensional families of probability distributions that remain stable under conditioning, and showing that this principle extends to mixtures and transformations, we demonstrate that a broader class of distributions can benefit from analytical conditioning. In short, analytical conditioning of multivariate densities can be extended far beyond the Gaussian case.
@@
This insight leads to a flexible generative method for estimating conditional distributions from data. The objective is to estimate the conditional distribution of a random vector $\mathbf{X_{1}}\in\mathbb{R}^{q}$ given an observation $\mathbf{x_{2}}$ of $\mathbf{X_{2}}\in\mathbb{R}^{p}$, based on a sample of moderate size and taking values in a continuous range, $(\mathbf{x_{1i}},\mathbf{x_{2i}})_{i=1}^{n}$ previously acquired. This setting is common in statistics, such as in probabilistic forecasting or imputation of systematically missing values, for example in medical or epidemiological applications.
Assuming conditions of existence to be fulfilled, a natural approach to estimate the conditional density $f(\mathbf{x_{1}}|\mathbf{x_{2}})$ is to first estimate the joint density by $\hat{f}(\mathbf{x_{1}},\mathbf{x_{2}})$ and then perform conditioning. Using our theoretical results, we show that this strategy can be applied far beyond the Gaussian case, which is often limiting due to its inability to capture complex dependencies and asymmetries typical in observed data. Notably, our approach allows modeling separately the dependence structure using copulas and the marginal distributions. Our method involves two main steps. First, we fit a copula to the data using a two-step procedure.
Next, we transform the original samples into a latent space, where conditioning is performed using analytical formulas and then we convert the samples back to the original space.
For applications, we will mainly focus on the Gaussian mixture copula model (GMCM) [tewari2011parametric] due to its ease of implementation and its ability to capture multi-modal dependencies. Additionally, we demonstrate that incorporating automatic differentiation enhances the parameter fitting process of GMCM compared to conventional approaches for computing derivatives [tewari2023estimation]. We conduct numerical experiments using both simulated and real data to demonstrate the relevance of our approach, evaluating the results with (multivariate) scoring rules.
@@
As a motivating example, we first consider here the Wine data-set, which represents chemical analysis of wines grown in the same region in Italy. The analysis determined the quantities of 13 constituents found in each of the three types of wines. For illustrative purposes, we focus on $2$ constituents, the Alcohol level and Malic acids concentration. Figure 1 illustrates the approach proposed using these variables. The joint density appears here to be rather well modeled by a GMCM, the conditional distribution of which can be estimated and simulated from efficiently as detailed throughout the paper.
@@
Figure 1: Steps of our conditional density estimation approach: from a sample of the joint distribution (left), we estimate the joint density by a mixture copular model (middle) and for for a given new value of Malic acids (5.04, red dotted line), we estimate (or sample from) the conditional distribution of Alcohol level by using analytical formulas (right).
@@
The remainder of the paper is organized as follows. In Section 2, we introduce the properties of stability by marginalization and conditioning for families of multivariate distributions, provide some example and counter-examples families and show that these properties can be extended to mixtures and transformations. In Section 3, we present our method for estimating efficiently a conditional distribution by assuming a parametric form for the joint distribution and leveraging analytical formulas for conditioning. Section 4 focuses on the implementation of our algorithm for a particular case which is the GMCM. For the demonstration of the usefulness of our method, we perform numerical simulations and include real data examples in Section 5. Finally, in Section 6, we draw conclusions and outline areas for future research.
@@ 2.1.1 Definition
Let us first consider a fixed dimension $d\geq 1$ and a product of spaces $\prod_{k=1}^{d}\mathcal{X}_{k}$. The $\mathcal{X}_{k}$ are assumed to be measure spaces, typically $\mathbb{R}$ or intervals, each endowed with the Lebesgue measure or restrictions thereof. We consider absolutely continuous probability distributions on $\prod_{k=1}^{d}\mathcal{X}_{k}$ with positive probability density functions $f^{(d)}(\cdot;\theta)$, where $\theta$ denotes a parameter. In this framework, we work indifferently with probability density functions (pdfs) or cumulative distribution functions (cdfs).
@@
Let us now introduce sets of indices that will be used for marginal distributions. For a given dimension $d>1$ the sets of $\ell$ strictly ordered indices are:
@@
  • $\mathbf{i}=(i_{1},\ldots,i_{\ell})$ set of $\ell$ indices with $1\leq i_{1}<\ldots<i_{\ell}\leq d$, $1\leq\ell\leq d-1$,
@@
For $f$ a positive pdf on $\prod_{k=1}^{d}\mathcal{X}_{k}$, $1\leq\ell\leq d-1$, and $(\mathbf{i},\mathbf{j})\in I_{\ell,d-\ell}^{d}$, we define the function:
$f_{\mathbf{i}}:(x_{i_{1}},\ldots,x_{i_{\ell}})\in\prod_{k=1}^{\ell}\mathcal{X}_{i_{k}}\rightarrow\int_{\prod_{k=1}^{d-\ell}\mathcal{X}_{j_{k}}}f(x)dx_{j_{1}}\ldots dx_{j_{d-\ell}}$
which applies the marginalization of a $d$-dimensional density over $d$-$\ell$ variables. Let us note that since $\mathbf{j}$ has $d$-$\ell$ components, it is completely defined by $\mathbf{i}$.
To extend this operation to a trans-dimensional setting, we extend the marginalization operator as follows.
For $d\geq 1$ and $f^{(d)}(\cdot;\theta)\in\mathcal{F}^{(d)}$, we define :
@@
We say that the trans-dimensional family of probability distributions $\mathcal{F}$ is stable by marginalization if it satisfies $\text{Marg}(\mathcal{F})\subset\mathcal{F}$.
Let us now consider conditioning. Since the pdfs considered here are positive, all considered conditional distributions are well-defined. For $f$ a positive pdf on $\prod_{k=1}^{d}\mathcal{X}_{k}$, a pair of sets of strictly ordered indices $(\mathbf{i},\mathbf{j})\in I_{\ell,m}^{d}$ and observed components $(x_{j_{1}},\ldots,x_{j_{m}})$, we denote:
@@
In the same manner as above for marginalization, we define a conditioning operator taking a pdf in a parametric family as input and returning all possible conditional densities i.e for all pairs of sets of strictly ordered indices and every possible set of conditioning variables. For a function $f^{(d)}(\cdot;\theta)\in\mathcal{F}^{(d)}$, we have :
@@
We say that trans-dimensional family of probability distributions $\mathcal{F}$ is stable by conditioning if it satisfies $\text{Cond}(\mathcal{F})\subset\mathcal{F}$.
@@ 2.1.2 Examples
Stability by marginalization and conditioning can be demonstrated by showing that the marginal or conditional distributions remain within the same family of distributions, albeit in a lower dimension. Here, we will present several examples of distribution families that exhibit this property, as well as two counter-examples.
@@
    where $\Theta^{(d)}=\mathbb{R}^{d}\times\text{PD}_{d}$ where $\text{PD}_{d}$ is the set of all symmetric positive definite matrices on $\mathbb{R}^{d\times d}$. The trans-dimensional family of Gaussian distributions is defined by: $\mathcal{F}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}$ with $D=+\infty$. By properties of the multivariate Gaussian distribution, we know that associated marginal and conditional distributions remain Gaussian with parameters that can be computed analytically via well-known formulas. We have that $\text{Marg}(\mathcal{F})\subset\mathcal{F}$ and $\text{Cond}(\mathcal{F})\subset\mathcal{F}$, and thus the trans-dimensional family of Gaussian distributions is stable by marginalization and by conditioning.
  • Student $t$-distribution: Another example is the family of multivariate Student t-distributions. Unlike the multivariate Gaussian distribution, it can model tail dependencies and it includes an additional “degrees of freedom” parameter, which controls the shape of the distribution, with higher values leading to distributions that resemble the multivariate Gaussian, while lower values result in heavier tails.
@@
    The multivariate Student $t$-distribution is stable by marginalization and conditioning [ding2016conditional]; more details are included in the appendix.
  • Other examples:
    Other examples of families of multivariate distributions stable by marginalization and conditioning are the multivariate unified skew normal and skew Student $t$-distributions [azzalini2013skew]. The family of elliptically contoured distributions also enjoys these properties [cambanis1981theory]. However, as shown in [fang2018symmetric], conditioning an elliptical distribution can alter its generator, and a closed-form expression for the conditional distribution is generally not available. This limitation restricts its applicability in the remainder of this paper. On a different note, Dirichlet distributions are stable by conditioning, up to scalings ensuring that resulting measures are supported on the unit simplex (in appropriate dimensions) [fang2018symmetric].
  • Counter example 1: Student $t$-distribution with fixed degrees of freedom parameter.
    Let us now consider a simple counter-example of a family of multivariate distributions which is stable by marginalization but not stable by conditioning. For a given $d\geq 1$, consider the family of multivariate Student $t$-distributions with a fixed degree of freedom $\nu\in\left[0,+\infty\right[$ i.e
@@
    where $f^{(d)}(\mathbf{x};\mu,\Sigma)=\frac{\Gamma(\frac{\nu+d}{2})}{\Gamma(\frac{\nu}{2})(\nu\pi)^{\frac{d}{2}}|\Sigma|^{\frac{1}{2}}}\left(1+\nu^{-1}(\mathbf{x}-\mu)^{T}\Sigma^{-1}(\mathbf{x}-\mu)\right)^{-\frac{(\nu+d)}{2}}$ and the trans-dimensional family is defined by $\mathcal{F}_{\nu}=\bigcup\limits_{d=1}^{D}\mathcal{F}^{(d)}_{\nu}$ with $D=+\infty$. Then, by properties of the multivariate Student $t$-distribution [ding2016conditional], this family is stable by marginalization. However, it is not stable by conditioning, because the resulting Student conditional distributions possess degrees of freedom different from $\nu$.
  • Counter example 2: $q$-Exponential distributions.
    A multivariate $q$-exponential distribution [li2023bayesian], denoted as $q\text{-ED}_{d}(\boldsymbol{\mu},\mathbf{C})$ for $(\boldsymbol{\mu},\mathbf{C})\in\mathbb{R}^{d}\times\text{PD}_{d}$, has the following density for $\mathbf{x}\in\mathbb{R}^{d}$ ($d\geq 1$ being fixed, here):
    $p(\mathbf{x}\mid\boldsymbol{\mu},\mathbf{C},q)=\frac{q}{2}(2\pi)^{-\frac{d}{2}}|\mathbf{C}|^{-\frac{1}{2}}r^{\frac{q}{2}-1}\exp\left(-\frac{r^{\frac{q}{2}}}{2}\right),\quad r(\mathbf{x})=(\mathbf{x}-\boldsymbol{\mu})^{\top}\mathbf{C}^{-1}(\mathbf{x}-\boldsymbol{\mu})$
    The counter-example presented now shows that for a given index $q$, the corresponding family of distributions is not stable by conditioning. From [li2023bayesian], we know that the $q$-exponential distribution is consistent in the sense defined by Kano [kano1994consistency]. By considering, $d=2$, $\mu=(0,0)^{t}$, $C=I_{2}$ and $\mathbf{x}=(x_{1},x_{2})^{t}$, we then have for $x_{1}\in\mathbb{R}$ :
    $p(x_{1})\propto(x_{1}^{2})^{{\frac{q}{2}}-1}\exp{\left(-\frac{1}{2}x_{1}^{q}\right)}$
    By assuming $p(x_{1})\neq 0$, the conditional distribution can then be written:
    $p(x_{2}|x_{1})\propto\frac{p(x_{1},x_{2})}{p(x_{1})}\propto\frac{(x_{1}^{2}+x_{2}^{2})^{(\frac{q}{2}-1)}}{(x_{1}^{2})^{(\frac{q}{2}-1)}}\exp{\left(-\frac{(x_{1}^{2}+x_{2}^{2})^{\frac{q}{2}}}{2}+\frac{(x_{1}^{2})^{\frac{q}{2}}}{2}\right)}$
    By taking $q=4$ and focusing on the exponential term, we get that the exponential term is:
    $\exp{\left(-\frac{x_{2}^{4}-2x_{1}^{2}x_{2}^{2}}{2}\right)}$
    If the conditional distribution were $q$-exponential (with $q=4$) it would be possible to write the exponential term as $\exp(-\frac{r^{2}}{2})$ with $r$ a quadratic form in $x_{2}$, which is not the case.
    We conclude that the conditional distribution is not $q$-exponential with $q=4$.